# Secure Smart Contract with Control Flow Integrity

### v1

@@ Abstract.
Smart contracts power decentralized financial (DeFi) services but are vulnerable to complex security exploits that can lead to significant financial losses. Existing security measures often fail to adequately protect these contracts due to the composability of DeFi protocols and the increasing sophistication of attacks. Through a large-scale empirical study of historical transactions from the $30$ hacked DeFi protocols, we discovered that while benign transactions typically exhibit a limited number of unique control flows, in stark contrast, attack transactions consistently introduce novel, previously unobserved control flows. Building on these insights, we developed CrossGuard, a novel framework that enforces control flow integrity in real-time to secure smart contracts. Crucially, CrossGuard does not require prior knowledge of specific hacks; instead, it dynamically enforces control flow whitelisting policies and applies simplification heuristics at runtime. This approach monitors and prevents potential attacks by reverting all transactions that do not adhere to the established control flow whitelisting rules. Our evaluation demonstrates that CrossGuard effectively blocks $28$ of the $30$ analyzed attacks when configured only once prior to contract deployment, maintaining a low false positive rate of $0.28\%$ and minimal additional gas costs. These results underscore the efficacy of applying control flow integrity to smart contracts, significantly enhancing security beyond traditional methods and addressing the evolving threat landscape in the DeFi ecosystem.
copyright: nonecopyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: none
@@ 1. Introduction
Blockchain technology has revolutionized the creation of global, secure, and programmable ledgers, fundamentally altering how digital transactions are conducted. Central to this innovation are smart contracts, which operate on blockchains, allowing developers to define and enforce complex transactional rules directly on the ledger. This capability has positioned blockchains and smart contracts as the backbone of various decentralized financial (DeFi) services. As of March 13th, 2025, the total value locked in 3,973 DeFi protocols has surged to approximately \$87.82 billion (DefiLlama, 2025a), highlighting the substantial economic impact and growth of this technology.
@@
However, the increasing reliance on smart contracts has brought security concerns to the forefront. By the same date, vulnerabilities in DeFi smart contracts have resulted in financial losses exceeding \$11.21 billion USD (DefiLlama, 2025b). In response to these growing threats, researchers have developed an array of program analysis and verification techniques and tools aimed at securing smart contracts (Luu et al., 2016; Torres et al., 2018; Liu et al., 2018; Feist et al., 2019; Zhang et al., 2019; Bose et al., 2022; Ma et al., 2021; Zheng et al., 2022; Wang et al., 2024; Albert et al., 2020; Zhang et al., 2023). Additionally, to mitigate risks and ensure the integrity of smart contracts, developers often commission thorough security audits to identify and address potential errors and vulnerabilities prior to the contract deployment.
@@
Despite these advancements in security measures, the evolving landscape of smart contracts has continually outpaced traditional defenses. Modern smart contracts are now designed with the flexibility to support the layering of additional contracts, a feature particularly vital in the decentralized finance (DeFi) ecosystem. In DeFi, smart contracts facilitate a diverse array of financial products and services, including lending, borrowing, trading, and yield farming. These interlinked contracts are often referred to as “DeFi legos,” emphasizing their modularity. The ability to combine various DeFi smart contracts, a concept known as “DeFi composability,” is widely regarded as one of the key advantages of DeFi (Popescu, 2020; Amler et al., 2021; Schär, 2021). However, this complexity and interdependence introduce significant challenges to securing smart contracts with conventional methods. The security of a DeFi protocol depends not only on the correct design and implementation of its own contracts but also on the integrity of external contracts it interacts with. Additionally, experienced users or attackers can deploy their own smart contracts to invoke functions across multiple DeFi protocols in arbitrary sequences. Considering all possible interactions a DeFi protocol may encounter prior to deployment is often infeasible for traditional security approaches.
@@
A critical observation in hack transaction analysis is that they often exploit unintended control flows across multiple functions, deviating from the original design intentions of the developers. For instance, re-entrancy attacks leverage an unforeseen recursive control flow through default handlers in custom contracts, enabling repeated execution of a critical function within one transaction. Similarly, flash loan attacks manipulate multiple functions in a precisely timed sequence, utilizing large asset transfers to coerce the victim contract into executing unfavorable trades.
To further explore this phenomenon, we conducted an empirical study analyzing in total $1,327,925$ historical transactions of $30$ compromised protocols on Ethereum. Our findings reveal that the number of different control flows is relatively constrained. In all but two cases, attack transactions introduced novel control flows that had never been observed in any previous transaction.
@@
Note that control flow integrity is a well-established security technique in traditional software to prevent memory attacks from hijacking a program’s control flow (Burow et al., 2017; Abadi et al., 2009; Göktas et al., 2014). Though our proposed techniques in this paper are dramatically different from prior control flow integrity techniques for traditional software, we adapt this concept here to emphasize the critical role of control flow in our approach.
A key challenge for CrossGuard arises from its fundamental design to only whitelist control flows rather than using prior knowledge of hacks to blacklist potentially malicious ones. This approach, while enhancing security by adhering strictly to known safe paths, could inherently lead to an increase in false positives if not meticulously managed. Although the number of unique control flows is inherently limited, a naive whitelisting approach could still produce numerous false positives or demand substantial human intervention. To address this issue, CrossGuard employs a set of heuristics designed to simplify the collected control flows. Specifically, CrossGuard excludes all read-only function calls from the control flows, as these calls do not alter the blockchain ledger’s state. Additionally, CrossGuard records the ledger state accessed by each function call and tracks the read-after-write dependencies of these calls. If function calls in a control flow trace do not have any dependencies with each other, CrossGuard will treat them as separate control flows and assess them with whitelisting policies individually. These heuristics effectively reduce redundant traces and significantly lower the false positive rate for CrossGuard.
Experimental Results: We evaluated CrossGuard on the deployed smart contracts and their transactions of the $30$ DeFi protocols included in our empirical study. Our results indicate that, when configured only once before deployment, CrossGuard can effectively prevent $28$ out of $30$ attacks analyzed in our study, maintaining a low average false positive rate of just 0.28%. Unlike traditional methods, CrossGuard does not depend on historical transactions. Despite this, CrossGuard still surpasses the state-of-the-art which instruments the smart contracts with invariants learned from historical transactions. Moreover, after implementing two optimization techniques, CrossGuard achieves a minimal gas consumption overhead of 15.52% on average. These results demonstrate the usefulness of our empirical findings and CrossGuard.
@@
2. Background
We now give a basic background of notions that later sections assume readers are familiar with. In particular, a blockchain is a decentralized and distributed ledger that records transactions across multiple machines. It is constituted of a sequence of blocks that are cryptographically linked together, and each blockchain contains a set of transactions between accounts on the blockchain. A transaction refers to the act of transferring assets or information on the blockchain. In the context of Ethereum blockchain, a transaction is public and it typically involves either sending Ether (the native cryptocurrency) or an Externally Owned Accounts (EOAs) invoking or deploying a smart contract. An EOA is controlled by an individual or entity through private keys without any associated code running on the blockchain.
Smart Contract is a self-executing piece of code deployed on the blockchain that automatically enforces and executes the terms of an agreement when predefined conditions are met. Accounts that contain smart contracts are called Contract Accounts. Ethereum Virtual Machine (EVM) is the runtime environment for executing smart contracts on the Ethereum blockchain. It acts as a decentralized computing machine that processes instructions and maintains the state of all smart contracts and accounts on the network. In Ethereum, gas is a unit that measures the computational effort required to execute transactions or smart contracts. Different opcodes, have varying gas costs, with opcodes involving storage read and write being particularly expensive due to their impact on the blockchain’s state. Recently, Ethereum introduced the EVM opcodes TLOAD and TSTORE in EIP-1153 (Ethereum Improvement Proposals, 2023) that were implemented in the Cancun upgrade (Kraken Exchange, 2024) in March 2024. Similar to the standard storage read and write, TLOAD and TSTORE handle temporary data within a single transaction and reset with each new transaction. These opcodes significantly reduce gas costs for runtime validation, offering a valuable opportunity to enhance control flow integrity in smart contracts at runtime.
DeFi Protocol is a decentralized financial system consisting of interconnected smart contracts to provide services like lending, borrowing, and trading without traditional intermediaries. ERC20 is a standard template for smart contracts specifically designed for creating fungible tokens on Ethereum. The majority of valuable tokens on Ethereum are built using this standard. Etherscan is the most widely-used blockchain explorer for Ethereum. Etherscan provides labels for addresses, identifying them as hackers, MEV bots, or protocol deployers. (Eth, 2024). Control Flow in software engineering is the sequence of instructions or statements executed in a program. In the context of smart contracts, we use control flow to denote the order of function invocations within a given protocol’s contracts during a single transaction.
3. Empirical Study - Understanding Control Flows in Hacked DeFi Protocols
In this section, we present an empirical study of control flows of historical transactions of victim DeFi protocols leading up to a hack. Typically, a victim protocol operates for a period before being compromised. During this operational phase, various blockchain actors may build upon the protocol, introducing novel control flows. Our goal is to thoroughly examine these control flows and answer the following research question:
@@ 3. Empirical Study - Understanding Control Flows in Hacked DeFi Protocols
RQ1: How do control flows in hack transactions differ from those in other (benign) transactions prior to a hack?
@@
To answer RQ1, we conducted a study on a systematically collected benchmark comprising 30 victim protocols involved in security hacking incidents. See Section 6.1 for detailed information about our benchmark collection methodology. For each hacking incident, we examine the nature and uniqueness of the control flows in the hack transactions, determining whether they differ from all previously observed transactions. The detailed study results are available in the “RQ1” column in Table 2 in Section 6.
@@
Results: Out of the $30$ studied hack incidents, $24$ demonstrated control flows that were distinct from any previously observed transaction patterns (marked as ✓in Table 2), highlighting the novel mechanisms by which these exploits were conducted. However, in $6$ cases — specifically involving the protocols bZx, VisorFi, UmbrellaNetwork, Opyn, DODO and Bedrock_DeFi — the control flows had been observed previously in benign transactions. A detailed investigation into these exceptions revealed insightful nuances. In particular, the UmbrellaNetwork hack (marked as ✗) was traced back to an integer underflow vulnerability. The Bedrock_DeFi hack (marked as ✗) was traced back to an issue which mistakenly set the conversion ratio of ETH/uniBTC to 1:1. These two exploits were executed through one function call to steal funds, which, while not novel in terms of control flow pattern, leveraged a specific code vulnerability to breach the protocol (QuillAudits Team, 2025; UNO Re, 2021). Catching these two hacks involves combining control flow analysis with data flow analysis.
The remaining 4 hacks presented a different complexity; each involved various types of re-entrancy attacks (marked as ). These attacks, while operating under the guise of familiar top-level function calls, manifested unique and sophisticated control flows at deeper interaction levels, which are not included in the simplified top-level control flow that we extract as described in Section 6.1. This shows that even though the simple control flow-based analysis described previously can detect $24$ hacks, it misses $4$ hacks that a more refined control flow-based analysis approach may help identify. Thus, we propose to enhance our control flow analysis framework to more accurately detect and classify hacks with intricate control flows such as the above $4$ re-entrancy hacks. The specifics of this improved approach will be elaborated upon in Section 5, aiming to better capture the subtleties of complex exploit strategies and refine our understanding of hack dynamics within DeFi protocols.
@@
Finding 1: Our findings indicate that the vast majority of hack transactions introduce unique control flows that differ from all previously observed benign transactions.
4. Preliminary and Definitions
[…]
Figure 1. Call Tree.
@@ 4. Preliminary and Definitions
We now present the formalization of the control flow of a transaction which our control flow integrity technique relies on. We use protected protocol, denoted as $P$, to refer to the set of smart contracts protected by our technique. We use $C_{i}$ to represent the $i^{th}$ protected contract, where $P=\{C_{1},C_{2},\dots,C_{j}\}$. Typically, $P$ consists of all the core contracts of the protocol, as specified by the developers of the protocol. We use external contracts, denoted as $E$, to refer to smart contracts that a protected protocol is built on top of but are external to the protected protocol, such as stablecoins or oracles. In addition, any contract that is neither part of the protected protocol nor considered external is referred to as an untrusted contract (denoted as $U$).
The call tree of a transaction $tx$, denoted as $CT(tx)$, represents the tree structure of function calls made during the execution of the transaction. Each node in the tree corresponds to a function call, and each directed edge represents the call dependency between two functions. An invocation from a transaction $tx$ with respect to a protected protocol $P$, denoted as $\iota(tx,P)$, is a tuple that corresponds to a sequence of function calls starting from an entry point (a function call) associated with one of the contracts in $P$. An invocation is considered simple if it does not involve any re-entrancy to the protected protocol, i.e., between the start and end of a function call we do not have a second call to the same function. Otherwise, it is considered re-entrant. The control flow of an invocation, denoted as $CF(\iota(tx,P))$ (in short $CF(\iota)$), consists of the start and end of a function call within a protected contract, abstracting away calls between protected or external contracts that occur during the invocation. If an untrusted contract is called between two calls to protected contracts (i.e., re-entrancy to $P$), both calls to the protected contracts are included in the control flow. The control flow of a transaction $tx$ w.r.t. a protected protocol $P$, $CF(tx,P)$ is defined as a sequence of control flows of invocations. Formally, $CF(tx,P)=\langle CF(\iota_{1}),CF(\iota_{2}),...\rangle$.
[…]

### v2

# Enforcing Control Flow Integrity on DeFi Smart Contracts

@@ Abstract.
Smart contracts power decentralized financial (DeFi) services but are vulnerable to security exploits that can lead to significant financial losses. Existing security measures often fail to adequately protect these contracts due to the composability of DeFi protocols and the increasing sophistication of attacks. Through a large-scale empirical study of historical transactions from the $37$ hacked DeFi protocols, we discovered that while benign transactions typically exhibit a limited number of unique control flows, in stark contrast, attack transactions consistently introduce novel, previously unobserved control flows. Building on these insights, we developed CrossGuard, a novel framework that enforces control flow integrity onchain to secure smart contracts. Crucially, CrossGuard does not require prior knowledge of specific hacks. Instead, configured only once at deployment, it enforces control flow whitelisting policies and applies simplification heuristics at runtime. This approach monitors and prevents potential attacks by reverting all transactions that do not adhere to the established control flow whitelisting rules. Our evaluation demonstrates that CrossGuard effectively blocks $35$ of the $37$ analyzed attacks when configured only once at contract deployment, maintaining a low false positive rate of $0.26\%$ and minimal additional gas costs. These results underscore the efficacy of applying control flow integrity to smart contracts, significantly enhancing security beyond traditional methods and addressing the evolving threat landscape in the DeFi ecosystem.
runtime validation, control flow integrity, dynamic analysis
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773266isbn: 979-8-4007-2025-3/2026/04ccs: Security and privacy Software security engineeringccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Blockchain technology has revolutionized the creation of global, secure, and programmable ledgers, fundamentally altering how digital transactions are conducted. Central to this innovation are smart contracts, which operate on blockchains, allowing developers to define and enforce complex transactional rules directly on the ledger. This capability has positioned smart contracts as the backbone of various decentralized financial (DeFi) services. As of March 13th, 2025, the total value locked in 3,973 DeFi protocols has surged to approximately \$87.82 billion (DefiLlama, 2025b), highlighting the substantial economic impact and growth of this technology.
@@
However, by the same date, vulnerabilities in DeFi smart contracts have resulted in financial losses exceeding \$11.21 billion USD (DefiLlama, 2025a). In response, researchers have developed program analysis and verification techniques to secure smart contracts (Luu et al., 2016; Torres et al., 2018; Liu et al., 2018; Feist et al., 2019; Zhang et al., 2019; Bose et al., 2022; Ma et al., 2021; Zheng et al., 2022; Wang et al., 2024; Albert et al., 2020; Zhang et al., 2023), and developers often commission security audits prior to deployment.
@@
Despite these advancements in security measures, the evolving landscape of smart contracts has continually outpaced traditional defenses. Modern smart contracts are now designed with the flexibility to support the layering of additional contracts, a feature particularly vital in DeFi ecosystem. In DeFi, smart contracts facilitate a diverse array of financial products and services, such as lending and yield farming. These interlinked contracts are often referred to as “DeFi legos,” emphasizing their modularity. The ability to combine various DeFi smart contracts, a concept known as “DeFi composability,” is widely regarded as one of the key advantages of DeFi (Popescu, 2020; Amler et al., 2021; Schär, 2021). However, this complexity and interdependence introduce significant challenges to securing smart contracts with conventional methods. The security of a DeFi protocol depends not only on the correct design and implementation of its own contracts but also on the integrity of external contracts it interacts with. Additionally, experienced users or attackers can deploy their own smart contracts to invoke functions across multiple DeFi protocols in arbitrary sequences. Considering all possible interactions a DeFi protocol may encounter prior to deployment is infeasible for traditional security approaches.
@@
A critical observation in hack transaction analysis is that they often exploit unintended control flows across multiple functions, deviating from the original design intentions of the developers. For instance, re-entrancy attacks leverage an unforeseen recursive control flow through default handlers in custom contracts, enabling repeated execution of a critical function within one transaction. Similarly, flash loan attacks manipulate multiple functions in a precisely timed sequence, utilizing large asset transfers to coerce the victim contract into executing unfavorable trades. We conducted an empirical study analyzing transaction histories of $37$ compromised Ethereum protocols. Our findings reveal that control flows are relatively constrained, with attack transactions introducing novel flows never observed previously in all but one hacked protocol.
@@
A key challenge arises from CrossGuard’s whitelisting-only design. This approach, while enhancing security by adhering strictly to known safe paths, could inherently lead to an increase in false positives if not meticulously managed. Although the number of unique control flows is inherently limited, a naive whitelisting approach could still produce numerous false positives or demand substantial human intervention. To address this issue, CrossGuard employs heuristics to simplify control flows: excluding read-only calls that don’t alter state, tracking read-after-write dependencies, and treating independent calls as separate flows. These heuristics effectively simplify control flows, and lower false positive rates.
Experimental Results: We evaluated CrossGuard on the deployed smart contracts and their transactions of the $37$ hacked DeFi protocols included in our empirical study. Our results indicate that, when configured only once at deployment, CrossGuard can effectively prevent $35$ out of $37$ attacks analyzed in our study, maintaining a low average false positive rate of just 0.26%. Unlike traditional methods, CrossGuard does not depend on historical transactions. Despite this, CrossGuard still surpasses the state-of-the-art which instruments the smart contracts with invariants learned from historical transactions. Moreover, after implementing four optimization techniques, CrossGuard achieves a minimal gas consumption overhead of 3.53% on average. These results demonstrate the usefulness of our empirical findings and CrossGuard.
@@
2. Background and Empirical Study
The Ethereum Virtual Machine (EVM) executes smart contracts and maintains network state. Gas measures computational effort for transaction execution, with storage operations being particularly expensive. The recent TLOAD and TSTORE opcodes (Ethereum Improvement Proposals, 2023; Kraken Exchange, 2024) provide temporary transaction-scoped storage at reduced gas costs, with variables automatically initialized to zero at transaction start. DeFi protocols are decentralized financial systems using interconnected smart contracts for lending, borrowing, and trading. Control Flow in software engineering is the order in which program instructions execute. In smart contracts, runtime control-flow analysis can be performed at different granularities, from low-level opcodes to high-level function calls, raising the question of which level offers the best trade-off between security coverage and computational cost.
Figure 1. Running example showing CrossGuard’s response to simple invocation (left) and re-entrancy attack (right)
Two-panel running example. Left: a benign/simple invocation sequence. Right: a re-entrancy attack sequence. The figure illustrates how CrossGuard observes function-level control flow and reacts differently in benign versus attack cases.
To ground our framework, we hypothesize that function-level control flows are sufficient to distinguish malicious from benign transactions, while remaining practical overheads for runtime monitoring. To evaluate this hypothesis, we analyze each victim protocol’s transaction history prior to the incident. Because protocols typically operate for some time before being exploited and attract many interacting actors, benign behavior can evolve over time, diversifying and introducing previously unseen control flows. This motivates the following research question:
@@ 2. Background and Empirical Study
RQ1: How do (function level) control flows in hack transactions differ from those in other (benign) transactions prior to a hack?
@@
To answer RQ1, we conducted a study on a systematically collected benchmark comprising $37$ victim protocols involved in security hacking incidents. The scope of this work is described in Section 4, and the details of our benchmark collection methodology are provided in Section 6.1. For each hacking incident, we examine the uniqueness of the control flows in the hack transactions, determining whether they differ from all previously observed transactions. The detailed study results are available in the “RQ1” column in Table 2 in Section 6. In this study, we only collect the top-level function calls to contracts of the victim protocol, ignoring deeper control flows within external calls.
@@
Results: Out of the $37$ studied hack incidents, $32$ demonstrated control flows that were distinct from any previously observed transaction patterns (marked as ✓in Table 2), highlighting the novel mechanisms by which these exploits were conducted. However, in $5$ cases (bZx, VisorFi, Opyn, DODO and Bedrock_DeFi), the control flows had been observed previously in benign transactions. A detailed investigation into these exceptions revealed insightful nuances. In particular, the Bedrock_DeFi hack (marked as ✗) was traced back to an issue which mistakenly set the conversion ratio of ETH/uniBTC to 1:1. This exploit was executed through one function call to steal funds, which, while not novel in terms of control flow pattern, leveraged a specific code vulnerability to breach the protocol (QuillAudits Team, 2025). Catching this hack involves combining control flow analysis with data flow analysis. The remaining $4$ hacks posed a different complexity: each involved various types of re-entrancy (marked as ). Although these attacks appeared to have familiar top-level function calls, they triggered unique and sophisticated control flows at deeper interaction levels. This shows that, even the simple control flow-based analysis can detect $32$ hacks, it misses $4$ that require a more refined control flow-based approach to identify. Thus, we propose to enhance the control flow analysis framework to more accurately detect and classify attacks with intricate control flows, such as these $4$ re-entrancy hacks. The details of this enhanced approach are presented in Section 5. Unless otherwise stated, we use the term “control flow” throughout the rest of this paper to refer to function-level control flow.
@@
Finding: the vast majority of hack transactions introduce unique function-level control flows that differ from all previously observed benign transactions.
3. Running Example
To illustrate how CrossGuard monitors and enforces control flow integrity, we present two contrasting scenarios: a benign transaction with simple invocation and a malicious re-entrancy attack. Figure 1 demonstrates CrossGuard’s response to these two different transactions. The left side of Figure 1 illustrates a benign transaction involving two protected contracts. Contract 1 contains FuncA, whose code externally invokes FuncB in Contract 2. When a user initiates this interaction, because they are not part of the protected protocol, they can only access FuncA_Untrusted(), which triggers the monitoring mechanism. Contract 1 first calls EnterFunc(A) to notify the guard contract of function entry, receiving an invocation number invNum=1 that uniquely identifies this execution count. Since FuncB is an expected external call within FuncA, and both contracts belong to the protected protocol, Contract 1 invokes FuncB_Trusted(invNum) directly, bypassing additional guard contract interactions while still passing the invocation number for runtime property tracking. Contract 2 monitors its execution properties and returns them to Contract 1, which then calls ExitFunc(A, properties) to report completion along with collected runtime properties. The guard contract ultimately records a simple control flow pattern <A-start, A-end>, which represents normal protocol behavior and is classified as benign.
The right side of Figure 1 demonstrates how CrossGuard stops a cross function re-entrancy attack. An attacker deploys Contract 4 and exploits Contract 3, which contains FuncC and FuncD that together constitute a re-entrancy vulnerability. The attack begins when the attacker invokes FuncC_Untrusted(), which subsequently triggers EnterFunc(C) and receiving invNum=1. During execution, Contract 3 invokes Attack() in the attacker’s Contract 4, which then re-enters Contract 3 and calls FuncD_Untrusted(). This triggers another EnterFunc(D) call, which returns the same invNum=1, indicating that this function call occurs within the context of the ongoing invocation. When both functions complete their execution, they call their respective ExitFunc methods, resulting in the guard contract recording a nested control flow pattern <C-start, <D-start, D-end>, C-end>. This nested structure clearly indicates a same-contract cross function re-entrancy behavior, and unless explicitly whitelisted by protocol administrators, CrossGuard blocks such transactions by default.
The example highlights several key aspects of CrossGuard’s design. The dual function variants(defined in Section 4) approach optimizes performance by allowing trusted intra-protocol calls to bypass guard contract interactions and save gas. The invocation numbering system enables correlation of related function calls within complex control flow sequences. Critically, both function variants collect runtime properties that are passed to the guard contract, providing additional context beyond control flow patterns to reduce false positives. The guard contract uses both control flow structure and runtime properties to distinguish between legitimate sequential execution patterns and malicious anomalies.
4. Definitions, Threat Model and Scope
@@
We now formalize the control flow of a transaction which CrossGuard relies on. We use protected protocol, denoted as $P$, to refer to the set of smart contracts protected by our technique. We use $C_{i}$ to represent the $i^{th}$ protected contract, where $P=\{C_{1},C_{2},\dots,C_{j}\}$. Typically, $P$ consists of all the core contracts of the protocol, as specified by the developers of the protocol. We use external contracts, denoted as $E$, to refer to smart contracts that a protected protocol is built on top of but are external to the protected protocol, such as stable coins or oracles. In addition, any contract that is neither part of the protected protocol nor considered external is referred to as an untrusted contract (denoted as $U$). All functions within protected contracts are referred to as protected functions.
Definition 4.0 (Tracked Runtime Properties).
As shown in Section 3, CrossGuard tracks runtime properties for each execution of a protected function, with the detailed algorithms given in Section 5.4. Specifically, it tracks the following two properties:
  • Runtime Read-Only ($\mathit{isRR}$): True if the function does not change any blockchain state on the fly.
  • Read-After-Write Dependency ($\mathit{isRAW}$): True if, within the same invocation count (as tracked by $\mathit{invNum}$), the function reads from any storage location that was written by a previous invocation in the same transaction.
Both properties are maintained as boolean flags and are mergeable under intra-protocol composition: for a protected function $f$ calling another protected function $g$ within the same invocation, $f$ aggregates $g$’s runtime properties upon return to reflect the combined behavior. The merged properties at return are updated as $\mathit{isRR}_{f}\coloneqq\mathit{isRR}_{f}\land\mathit{isRR}_{g},\quad\mathit{isRAW}_{f}\coloneqq\mathit{isRAW}_{f}\lor\mathit{isRAW}_{g}.$
CrossGuard instruments each protected contract function with two variants to optimize monitoring overhead while maintaining security guarantees. For each function $f$ in a protected contract $C_{i}\in P$, CrossGuard creates two variants. Both variants collect and propagates runtime properties.
  • Untrusted variant $f\_\text{Untrusted}$: Accessible to any caller but mandatory triggers EnterFunc and ExitFunc of guard contract.
  • Trusted variant $f\_\text{Trusted}$: Restricted to calls from other protected contracts via access control, bypasses guard contract interactions.
Definition 4.0 (Call Tree and Invocation).
The call tree $CT(tx)$ of transaction $tx$ represents the tree structure of function calls, where nodes correspond to function calls and directed edges represent call dependencies. An invocation $\iota(tx,P)$ is a sequence of function calls starting from an entry point in protected protocol $P$. Invocations are classified as re-entrant (containing recursive calls to protected contracts through untrusted intermediaries), or simple (no re-entrancy to protected contracts).
Definition 4.0 (Control Flow).
The control flow of an invocation $CF(\iota)$ captures the start and end of function calls within protected contracts, abstracting away calls to external contracts unless they facilitate re-entrancy to protected contracts. The control flow of a transaction $CF(tx,P)$ is defined as the sequence of all invocation control flows $\langle CF(\iota_{1}),CF(\iota_{2}),\ldots,CF(\iota_{n})\rangle$. A control flow is trivial only if it consists of a simple invocation, otherwise it is non-trivial.
Consider the re-entrancy scenario in Section 3. The two invocations would be $\iota_{1}$ (the initial FuncC call) and $\iota_{2}$ (the re-entrant FuncD call). The resulting control flow $CF(tx,P)=\langle\langle C_{3}.\texttt{C-s},\langle C_{3}.\texttt{D-s},\\ C_{3}.\texttt{D-e}\rangle,C_{3}.\texttt{C-e}\rangle\rangle$ clearly reveals re-entrancy, where notation ‘-s’ and ‘-e’ denote function start and end respectively.
Threat Model: Our threat model assumes sophisticated attackers who can deploy arbitrary untrusted contracts to interact with protected protocols. Given the transparent nature of blockchain systems, attackers have access to CrossGuard’s on-chain implementation, enabling them to probe its detection mechanisms and whitelisting policies/heuristics. Additionally, attackers may attempt to evade detection by decomposing complex attacks into multiple simpler transactions or adapting their strategies to minimize detection risks while maximizing profit. We also assume attackers do not have administrative access to the protected protocol.
Scope: CrossGuard blocks attacks that exhibit non-trivial control flows(those utilizing multiple function invocations), or re-entrancy patterns. CrossGuard cannot prevent vulnerabilities exploitable through single function calls with trivial control flows, such as access control bypasses or bridge exploits.